Selected operating data for two divisions of Outback Brewing, Ltd., of Australia are given below: Division Queensland New South Wales Sales $ 1,071,000 $ 2,016,000 Average operating assets $ 510,000 $ 560,000 Net operating income $ 107,100 $ 141,120 Property, plant, and equipment (net) $ 251,000 $ 201,000 Required: 1. Compute the rate of return for each division using the return on investment (ROI) formula stated in terms of margin and turnover. (Do not round intermediate calculations. Round your answers to 2 decimal places.) Queensland division%times%New South Wales division%times% Margin, Turnover, ROI for both
Answer #1

Queensland Division

Margin = Net operating income/Sales

= 107,100/1,071,000

= 10%

Assets turnover = Sales/Average operating assets

= 1,071,000/510,000

= 2.1

Return on investment = Margin x Assets turnover

= 10% x 2.1

= 21%

New South Wales Division

Margin = Net operating income/Sales

= 141,120/2,016,000

= 7%

Assets turnover = Sales/Average operating assets

= 2,016,000/560,000

= 3.6

Return on investment = Margin x Assets turnover

= 7% x 3.6

= 25.2%

Kindly comment if you need further assistance. Thanks